# Optimal Methods for Unknown Piecewise Smooth Problems I: Convex Optimization

### v1

@@ 6 Restarted APEX for unknown $\mu^{*}$
In this section, we propose an algorithm that eliminates the requirement for a priori knowledge of the QG constant $\mu^{*}$. To achieve this, we embed Algorithm 4 within a systematic ”guess–and–check” mechanism, yielding Algorithm 6. At each outer iteration, we maintain an estimate $\hat{\mu}_{s}>0$ of the true QG constant $\mu^{*}$ and use it to define a candidate optimality gap $\tilde{\Delta}_{s}$, which governs two verification mechanisms: $(1)$ Within a fixed budget of inner iterations, $\mathcal{AWG}$ attempts to produce a W-certificate for the current reference point $\bar{y}$. Failure to generate such a certificate implies that $\tilde{\Delta}_{s}$ underestimates the true optimality gap $f\left(\bar{y}^{s-1}\right)-f^{*}$, which, under the QG assumption, can only occur if $\hat{\mu}_{s}>\mu^{*}$. In that case (Line 6 of Algorithm 4), $\hat{\mu}_{s}$ is halved and the guess–and–check loop is restarted. $(2)$ If $\mathcal{AWG}$ succeeds, we invoke $\mathcal{AGR}$ to simultaneously refine the upper bound of $f\left(x\right)$ and the lower bound estimate. Similar to Algorithm 4, $\mathcal{AGR}$ will make progress (upper bound or lower bound) before exhausting its budget if $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\tilde{\Delta}_{s}$ holds. If instead, neither the upper bound nor the certificate-based lower bound improves within the allotted iterations, i.e., $\mathcal{AGR}$ enters Line 10. We again infer $\tilde{\Delta}_{s}$ was set too small (implies $\hat{\mu}_{s}>\mu^{*}$), and we halve $\hat{\mu}_{s}$ in Line 11 of Algorithm 6.
@@
4:  Compute $\bar{L}\left(t\right)$ (defined in (8)) based on $\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|$ and $\mathcal{P}=\mathcal{P}\cup\mathcal{P}_{t}$
5:  if $\left\|x^{t}-\bar{y}\right\|^{2}\geq\frac{2\theta\Delta}{\mu}$ then
6:    Return $\text{{True}},\hat{x}^{t},\text{{\text{False}}},\mathcal{P}\cap\mathcal{B}\left(\bar{y},\sqrt{\frac{2\theta\Delta}{\mu}}\right)$
@@
8:    Return $\text{{True}},\hat{x}^{t},\text{{True}},\ldots$
9:  else if $t\geq\left\lceil\sqrt{\frac{6\theta}{2\theta-1}+\frac{4\bar{L}\left(t\right)\theta}{\mu\left(2\theta-1\right)}}\right\rceil$ then
10:    Return $\text{{\text{False}}},\hat{x}^{t},\ldots$
11:  end if
12:end for
@@
2:Initialize: upper bound $\bar{f}=f\left(\bar{y}^{0}\right)$, $\tilde{\Delta}_{1}=\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}\ ,\ \underline{f}=\bar{f}-\tilde{\Delta}_{1}\ ,\ \mu_{0}=1\ ,\ \tilde{\Delta}_{0}=+\infty$
@@
4:   $\left(\text{{Half-Flag}},\hat{x}^{t},\tilde{\mathcal{P}}_{s},\ldots\right)\leftarrow\mathcal{AWG}\left(f\left(\cdot\right),\bar{y}^{s-1},\tilde{\Delta}_{s},m,\sqrt{\frac{2\left(1+\beta\right)\tilde{\Delta}_{s}}{\hat{\mu}_{s}}},\beta\right)$
5:  if Half-Flag is False then
@@
9:   $\left(\text{{Half-Flag}},\hat{x}^{t},\text{{Lower-Flag}},\tilde{\mathcal{P}}_{s}\right)\leftarrow\mathcal{AGR}\left(f\left(\cdot\right),\underline{f},\tilde{\Delta}_{s},\bar{y}^{s-1},\hat{\mu}_{s},\theta,m\right)$
10:  if Half-Flag is False then
11:    $\hat{\mu}_{s}=\hat{\mu}_{s}/4$, $\tilde{\Delta}_{s}=\min\left\{{4\Delta_{s}},\frac{2}{\hat{\mu}_{s}}\left\|f^{\prime}\left(\bar{y}^{s-1}\right)\right\|^{2},\frac{9\mu_{s-1}}{4\hat{\mu}_{s}}\Delta_{s-1}\right\},\underline{f}=\bar{f}-\tilde{\Delta}_{s}$ and go to Line 4
12:  else if Lower-Flag is False then
13:    $\tilde{\Delta}_{s}=\theta\tilde{\Delta}_{s},\underline{f}=\bar{f}-\tilde{\Delta}_{s},\Delta_{s}=\tilde{\Delta}_{s},\mu_{s}=\hat{\mu}_{s},\iota_{s}=\sqrt{\frac{2\Delta_{s}}{\mu_{s}}},\mathcal{P}_{s}=\tilde{\mathcal{P}}_{s}$ and go to Line 9
14:  end if
15:   $\bar{y}^{s}=\hat{x}^{t},\bar{f}=f\left(\bar{y}^{s}\right),\tilde{\Delta}_{s+1}=\bar{f}-\underline{f}$
16:end for
Beyond the “guess-and-check” procedure for validating $\hat{\mu}_{s}$, Algorithm 6 crucially relies on a certificate-transfer mechanism to carry forward valid W-certificate and thus maintain sharp lower bounds on $f^{*}$ even the QG modulus is misspecified. Concretely, suppose at the $s$-th outer stage $(s>1)$, Algorithm 6 enters Line 6, so that $\mathcal{AWG}$ fails to produce a new W-certificate for the current iterate $\bar{y}^{s-1}$. Although no fresh W-certificate for $\bar{y}^{s-1}$, we still hold that the most recent valid $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate that was generated for the previous iterate $\bar{y}^{s-2}$. In the following Lemma 6.1, we show that a W-certificate of $\bar{y}^{s-1}$ can be transferred from $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate of $\bar{y}^{s-2}$. By invoking this transferred certificate, we obtain a valid and typically much tighter–lower bound on $f^{*}$ than the coarse estimate $f\left(\bar{y}^{s-1}\right)-\left\|f^{\prime}\left(\bar{y}^{s-1}\right)\right\|^{2}\big/\left(2\hat{\mu}_{s}\right)$, thereby preserving the accelerated convergence rate. An analogous argument applies at the second halving step (Line 11). This strategy is for ensuring that all lower-bound estimates remain as tight as possible throughout the algorithm.
@@
$\begin{split}\tilde{\iota}\mathcal{V}_{\tilde{\mathcal{P}},f}\left(\tilde{\iota};\tilde{y}\right)&=\max_{x\in\mathcal{B}\left(\tilde{y};\tilde{\iota}\right)}\left\{\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(a)}{\leq}\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\max_{x\in\mathcal{B}\left(\bar{y};\left(1+2c\right)\iota\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &=\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\iota\mathcal{V}_{\mathcal{P},f}\left(\left(1+2c\right)\iota;\bar{y}\right)\\ &\overset{(b)}{\leq}\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\iota\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)\\ &=\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\max_{x\in\mathcal{B}\left(\tilde{y};\iota\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(c)}{\leq}f\left(\tilde{y}\right)+cf\left(\bar{y}\right)-\left(1+2c\right)\tilde{l}\end{split}$ (29)
Here, $(a)$ holds by taking the definition of $\tilde{\iota}$ and $\left\|\bar{y}-\tilde{y}\right\|\leq c\iota$, $(b)$ follows from $\mathcal{V}_{\mathcal{P},f}\left(\left(1+2c\right)\iota;\bar{y}\right)\leq\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ since it is a non-increasing function by Lemma 4.1 and $(c)$ holds by function $f$ is convex function and $\min_{x\in\mathcal{B}\left(\bar{y};\iota\right)\cap X}\psi_{\mathcal{P},f}\left(x\right)\geq\tilde{l}$. Hence, we complete our proof. ∎∎
We are now ready to present the main convergence theorem for Algorithm 6. Collectively, the preceding lemmas ensure that the lower bound estimate of $f^{*}$ remains valid under the current QG constant estimate $\hat{\mu}_{s}$, as established in Lemma 6.3. Building on this foundation, Theorem 6.1 provides the complete convergence guarantee.
@@
Algorithm 6 will generate W-certificate in Line 8 or Line 13. Specifically,
@@
    If Algorithm 6 reaches Line 13, then $\mathcal{AGR}$ will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. Here $\Delta_{s}$ is updated newly in Line 13.
@@
Part 1: It follows that $\mathcal{AWG}$ will return a W-certificate unless it enters Line 10 in Algorithm 3, i.e., it return Half-Flag equals to False. Hence, if Algorithm 6 runs to Line 8, the $\mathcal{AWG}$ subroutine has generated a W-certificate successfully. Specifically, by Proposition 4.1, the return point set $\mathcal{P}_{s}$ is an $\left(\iota_{\max},(1+\beta)\Delta/\iota_{\max}\right)=\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$.
Part 2: The proof process for Algorithm 6 enters Line 13 is similar to Proposition 4.1. ∎∎
@@
Suppose function $f$ is convex satisfies $\hat{\mu}_{s}$-QG, then Algorithm 6 updates $\underline{f}$ at Lines 6, 11 and 13 and the lower bound $\underline{f}$ satisfying $\underline{f}\leq\min_{x\in X}f\left(x\right)$.
@@
The parameter $\underline{f}$ is updated in three situations (Line 6, 11 and 13) within the algorithm. We will analyze each case to demonstrate that $\underline{f}$ is always set to a value less than or equal to the optimum $f^{*}$. The value $\tilde{\Delta}_{s}$ is chosen as the minimum of several estimates for the optimality gap $f\left(\bar{y}^{s-1}\right)-\min_{x\in X}f\left(x\right)$, so proving that all of these estimates are valid upper bounds on the gap is needed.
@@
$\begin{split}f^{*}&\geq f\left(x\right)+\left\langle f^{\prime}\left(x\right),x^{*}-x\right\rangle\geq f\left(x\right)-\left\|f^{\prime}\left(x\right)\right\|\left\|x^{*}-x\right\|\\ &\geq f\left(x\right)-\left\|f^{\prime}\left(x\right)\right\|\sqrt{2\left(f\left(x\right)-f^{*}\right)/\hat{\mu}_{s}},\end{split}\ ,$ (30)
which implies that $f^{*}\geq f\left(x\right)-2\left\|f^{\prime}\left(x\right)\right\|^{2}/\hat{\mu}_{s}$. A more refined bound can be derived by using the $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate, denote $\mathcal{P}_{s-1}$, from the previous outer iteration. This certificate at $\bar{y}^{s-2}$ can be transformed into a new certificate for point $\bar{y}^{s-1}$. Specifically, if Algorithm 6 enters Line 6 for $s\geq 2$, there exists a point set $\mathcal{P}_{s-1}$ as $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate for point $\bar{y}^{s-2}$ in function $f$ by Lemma 6.2. Since $\bar{y}^{s-1}$ is generated in Line 15 of Algorithm 6, it means the $\left\|\bar{y}^{s-1}-\bar{y}^{s-2}\right\|\leq\sqrt{2\theta\tilde{\Delta}_{s-1}/\mu_{s-1}}$, otherwise Algorithm 6 would enter Line 13. By the updating rule, we have $\theta\tilde{\Delta}_{s-1}\leq\Delta_{s-1}$, which implies that $\left\|\bar{y}^{s-1}-\bar{y}^{s-2}\right\|\leq\sqrt{2\Delta_{s-1}/\mu_{s-1}}=\iota_{s-1}$. Furthermore, by Lemma 4.3 indicating $\min_{x\in\mathcal{B}\left(\bar{y}^{s-2};\iota_{s-1}\right)}\psi_{\mathcal{P}_{s-1},f}\left(x\right)\geq f\left(\bar{y}^{s-2}\right)-\Delta_{s-1}$. Hence, applying Lemma 6.1 with $\mathcal{P}=\mathcal{P}_{s-1}$, $\iota=\iota_{s-1}$, $\nu=\nu_{s-1}$, $\bar{y}=\bar{y}^{s-2},$ $\tilde{l}=f\left(\bar{y}^{s-2}\right)-\Delta_{s-1},\tilde{y}=\bar{y}^{s-1}$, and $c=1$ yields that $\tilde{\mathcal{P}}_{s}=\left\{\bar{y}^{s-1}\right\}\cup\mathcal{P}_{s-1}$ is an $\left(\tilde{\iota}_{s},\tilde{\nu}_{s}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$, where $\tilde{\iota}_{s}=2\iota_{s-1}$ and
$\tilde{\nu}_{s}=\frac{1}{2\iota_{s-1}}\left(f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)+3\Delta_{s-1}\right)\ .$ (31)
@@
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\zeta_{s},\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2}\right\}\ .$ (32)
@@
$2\hat{\mu}_{s}\iota_{s-1}^{2}=\frac{4\hat{\mu}_{s}}{\mu_{s-1}}\Delta_{s-1}\overset{(a)}{\geq}\frac{4\hat{\mu}_{s}}{3\mu_{s-1}}\zeta_{s}\ ,$ (33)
where $\left(a\right)$ holds by $f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)\leq 0$. $\zeta_{s}>0$ and (33) imply $3\mu_{s-1}\zeta_{s}/4\hat{\mu}_{s}$ is an upper bound of the second term in (32). Note that $\hat{\mu}_{s}\leq\mu_{s-1}/2$, hence, we have
$\frac{3\mu_{s-1}}{4\hat{\mu}_{s}}\zeta_{s}\geq\max\left\{\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2},\frac{3}{2}\zeta_{s}\right\}\ .$ (34)
Hence, combining (34) and (32) yields
$\begin{split}f\left(\bar{y}^{s-1}\right)-f^{*}&\leq\max\left\{\zeta_{s},\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2},\frac{3}{2}\zeta_{s}\right\}\leq\frac{3\mu_{s-1}}{4\hat{\mu}_{s}}\zeta_{s}\end{split}\overset{(a)}{\leq}\frac{9\mu_{s-1}}{4\hat{\mu}_{s}}\Delta_{s-1}\ ,$ (35)

### v2

@@ 6 Restarted APEX for unknown $\mu^{*}$
In this section, we propose an algorithm that eliminates the requirement for a priori knowledge of the QG constant $\mu^{*}$. To achieve this, we embed Algorithm 4 within a systematic ”guess–and–check” mechanism, yielding Algorithm 6. At each outer iteration, we maintain an estimate $\hat{\mu}_{s}>0$ of the true QG constant $\mu^{*}$ and use it to define a candidate optimality gap $\tilde{\Delta}_{s}$, which governs two verification mechanisms: $(1)$ Within a fixed budget of inner iterations, $\mathcal{AWG}$ attempts to produce a W-certificate for the current reference point $\bar{y}$. Failure to generate such a certificate implies that $\tilde{\Delta}_{s}$ underestimates the true optimality gap $f\left(\bar{y}^{s-1}\right)-f^{*}$, which, under the QG assumption, can only occur if $\hat{\mu}_{s}>\mu^{*}$. In that case (Line 6 of Algorithm 4), $\hat{\mu}_{s}$ is halved and the guess–and–check loop is restarted. $(2)$ If $\mathcal{AWG}$ succeeds, we invoke $\mathcal{AGR}$ to simultaneously refine the upper bound of $f\left(x\right)$ and the lower bound estimate. Similar to Algorithm 4, $\mathcal{AGR}$ will make progress (upper bound or lower bound) before exhausting its budget if $f\left(\bar{y}^{s-1}\right)-f^{*}\leq\tilde{\Delta}_{s}$ holds (see Proposition 6.1 for more details).
@@
4:  Compute $\bar{L}\left(t\right)$ (defined in (3.2)) based on $\left\|x^{t,l_{t}}-x^{t,r_{t}}\right\|$ and $\mathcal{P}=\mathcal{P}\cup\mathcal{P}_{t}$
5:  if Flag is True or $\left\|x^{t}-\bar{y}\right\|^{2}\geq\frac{2\theta\Delta}{\mu}$ then
6:    Return $\hat{x}^{t},\text{{True}},\mathcal{P}\cap\mathcal{B}\left(\bar{y},\sqrt{\frac{2\theta\Delta}{\mu}}\right)$
@@
8:    Return $\hat{x}^{t},\text{{\text{False}}},\ldots$
9:  end if
10:end for
@@
2:Initialize: upper bound $\bar{f}=f\left(\bar{y}^{0}\right)$, $\tilde{\Delta}_{1}=\frac{2\left\|f^{\prime}\left(\bar{y}^{0}\right)\right\|^{2}}{\hat{\mu}_{1}}\ ,\ \underline{f}=\bar{f}-\tilde{\Delta}_{1}\ ,\ \mu_{0}=1\ ,\ {\Delta}_{0}=+\infty$
@@
4:   $\left(\text{{Quar-Flag}},\hat{x}^{s},\tilde{\mathcal{P}}_{s},\ldots\right)\leftarrow\mathcal{AWG}\left(f\left(\cdot\right),\bar{y}^{s-1},\tilde{\Delta}_{s},m,\sqrt{\frac{2\left(1+\beta\right)\tilde{\Delta}_{s}}{\hat{\mu}_{s}}},\beta\right)$
5:  if Quar-Flag is False then
@@
9:   $\left(\hat{x}^{s},\text{{Lower-Flag}},\tilde{\mathcal{P}}_{s}\right)\leftarrow\mathcal{AGR}\left(f\left(\cdot\right),\underline{f},\tilde{\Delta}_{s},\bar{y}^{s-1},\hat{\mu}_{s},\theta,m\right)$
10:  if Lower-Flag is True then
11:    $\tilde{\Delta}_{s}=\theta\tilde{\Delta}_{s},\underline{f}=\bar{f}-\tilde{\Delta}_{s},\Delta_{s}=\tilde{\Delta}_{s},\mu_{s}=\hat{\mu}_{s},\iota_{s}=\sqrt{\frac{2\Delta_{s}}{\mu_{s}}},\mathcal{P}_{s}=\tilde{\mathcal{P}}_{s}$ and go to Line 9
12:  end if
13:  if $f(\hat{x}^{s})<\underline{f}$ then
14:    go to Line 6
15:  end if
16:   $\bar{y}^{s}=\hat{x}^{s},\bar{f}=f\left(\bar{y}^{s}\right),\tilde{\Delta}_{s+1}=\bar{f}-\underline{f}$
17:end for
Beyond the “guess-and-check” procedure for validating $\hat{\mu}_{s}$, Algorithm 6 crucially relies on a certificate-transfer mechanism to carry forward valid W-certificate and thus maintain sharp lower bounds on $f^{*}$ even the QG modulus is misspecified. Concretely, suppose at the $s$-th outer stage $(s>1)$, Algorithm 6 enters Line 6, so that $\mathcal{AWG}$ fails to produce a new W-certificate for the current iterate $\bar{y}^{s-1}$. Although no fresh W-certificate for $\bar{y}^{s-1}$, we still hold that the most recent valid $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate that was generated for the previous iterate $\bar{y}^{s-2}$. In the following Lemma 6.1, we show that a W-certificate of $\bar{y}^{s-1}$ can be transferred from $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate of $\bar{y}^{s-2}$. By invoking this transferred certificate, we obtain a valid and typically much tighter–lower bound on $f^{*}$ than the coarse estimate $f\left(\bar{y}^{s-1}\right)-\left\|f^{\prime}\left(\bar{y}^{s-1}\right)\right\|^{2}\big/\left(2\hat{\mu}_{s}\right)$, thereby preserving the accelerated convergence rate.
This strategy is for ensuring that all lower-bound estimates remain as tight as possible throughout the algorithm.
@@
$\begin{split}\tilde{\iota}\mathcal{V}_{\tilde{\mathcal{P}},f}\left(\tilde{\iota};\tilde{y}\right)&=\max_{x\in\mathcal{B}\left(\tilde{y};\tilde{\iota}\right)}\left\{\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(a)}{\leq}\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\max_{x\in\mathcal{B}\left(\bar{y};\left(1+2c\right)\iota\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &=\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\iota\mathcal{V}_{\mathcal{P},f}\left(\left(1+2c\right)\iota;\bar{y}\right)\\ &\overset{(b)}{\leq}\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\iota\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)\\ &=\psi_{\mathcal{P},f}\left(\tilde{y}\right)-\psi_{\mathcal{P},f}\left(\bar{y}\right)+\left(1+2c\right)\max_{x\in\mathcal{B}\left(\tilde{y};\iota\right)}\left\{\psi_{\mathcal{P},f}\left(\bar{y}\right)-\psi_{\mathcal{P},f}\left(x\right)\right\}\\ &\overset{(c)}{\leq}f\left(\tilde{y}\right)+cf\left(\bar{y}\right)-\left(1+2c\right)\tilde{l}\end{split}$ (6.2)
Here, $(a)$ holds by taking the definition of $\tilde{\iota}$ and $\left\|\bar{y}-\tilde{y}\right\|\leq c\iota$, $(b)$ follows from $\mathcal{V}_{\mathcal{P},f}\left(\left(1+2c\right)\iota;\bar{y}\right)\leq\mathcal{V}_{\mathcal{P},f}\left(\iota;\bar{y}\right)$ since it is a non-increasing function by Lemma 4.1 and $(c)$ holds by function $f$ is convex function and $\min_{x\in\mathcal{B}\left(\bar{y};\iota\right)\cap X}\psi_{\mathcal{P},f}\left(x\right)\geq\tilde{l}$. Hence, we complete our proof. ∎
We are now ready to present the main convergence theorem for Algorithm 6. Collectively, the preceding lemmas (Lemma 6.2 and 6.3) ensure that the lower bound estimate of $f^{*}$ remains valid under the current QG constant estimate $\hat{\mu}_{s}$. Lemma 6.4 gives the result that the upper bound of total number to quarter estimated QG $\hat{\mu}_{s}$. Proposition 6.1 gives the complexity result for subroutine $\mathcal{AGR}$. Building on this foundation, Theorem 6.1 provides the complete convergence guarantee.
@@
Algorithm 6 will generate W-certificate in Line 8 or Line 11. Specifically,
@@
    If Algorithm 6 reaches Line 11, then $\mathcal{AGR}$ will generate a point set $\mathcal{P}_{s}$ as $\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$. Here $\Delta_{s}$ is updated newly in Line 11.
@@
Part 1: It follows that $\mathcal{AWG}$ will return a W-certificate unless it enters Line 10 in Algorithm 3, i.e., it return Quar-Flag equals to False. Hence, if Algorithm 6 runs to Line 8, the $\mathcal{AWG}$ subroutine has generated a W-certificate successfully. Specifically, by Proposition 4.1, the return point set $\mathcal{P}_{s}$ is an $\left(\iota_{\max},(1+\beta)\Delta/\iota_{\max}\right)=\left(\sqrt{2\Delta_{s}/\mu_{s}},\sqrt{\mu_{s}\Delta_{s}/2}\right)$-W-certificate for point $\bar{y}^{s-1}$.
Part 2: The proof process for Algorithm 6 enters Line 11 is similar to Proposition 4.1. ∎∎
@@
Suppose function $f$ is convex satisfies $\hat{\mu}_{s}$-QG, then Algorithm 6 updates $\underline{f}$ at Lines 6 and $1$ and the lower bound $\underline{f}$ satisfying $\underline{f}\leq\min_{x\in X}f\left(x\right)$.
@@
The parameter $\underline{f}$ is updated in two situations (Line 6 and 11) within the algorithm. We will analyze each case to demonstrate that $\underline{f}$ is always set to a value less than or equal to the optimum $f^{*}$. The value $\tilde{\Delta}_{s}$ is chosen as the minimum of several estimates for the optimality gap $f\left(\bar{y}^{s-1}\right)-\min_{x\in X}f\left(x\right)$, so proving that all of these estimates are valid upper bounds on the gap is needed.
@@
$\begin{split}f^{*}&\geq f\left(x\right)+\left\langle f^{\prime}\left(x\right),x^{*}-x\right\rangle\geq f\left(x\right)-\left\|f^{\prime}\left(x\right)\right\|\left\|x^{*}-x\right\|\\ &\geq f\left(x\right)-\left\|f^{\prime}\left(x\right)\right\|\sqrt{2\left(f\left(x\right)-f^{*}\right)/\hat{\mu}_{s}},\end{split}\ ,$ (6.3)
which implies that $f^{*}\geq f\left(x\right)-2\left\|f^{\prime}\left(x\right)\right\|^{2}/\hat{\mu}_{s}$. A more refined bound can be derived by using the $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate, denote $\mathcal{P}_{s-1}$, from the previous outer iteration. This certificate at $\bar{y}^{s-2}$ can be transformed into a new certificate for point $\bar{y}^{s-1}$. Specifically, if Algorithm 6 enters Line 6 for $s\geq 2$, there exists a point set $\mathcal{P}_{s-1}$ as $\left(\iota_{s-1},\nu_{s-1}\right)$-W-certificate for point $\bar{y}^{s-2}$ in function $f$ by Lemma 6.2. Since $\bar{y}^{s-1}$ is generated in Line 16 of Algorithm 6, it means the $\left\|\bar{y}^{s-1}-\bar{y}^{s-2}\right\|\leq\sqrt{2\theta\tilde{\Delta}_{s-1}/\mu_{s-1}}$, otherwise Algorithm 6 would enter Line 11. By the updating rule, we have $\theta\tilde{\Delta}_{s-1}\leq\Delta_{s-1}$, which implies that $\left\|\bar{y}^{s-1}-\bar{y}^{s-2}\right\|\leq\sqrt{2\Delta_{s-1}/\mu_{s-1}}=\iota_{s-1}$. Furthermore, by Lemma 4.3 indicating $\min_{x\in\mathcal{B}\left(\bar{y}^{s-2};\iota_{s-1}\right)}\psi_{\mathcal{P}_{s-1},f}\left(x\right)\geq f\left(\bar{y}^{s-2}\right)-\Delta_{s-1}$. Hence, applying Lemma 6.1 with $\mathcal{P}=\mathcal{P}_{s-1}$, $\iota=\iota_{s-1}$, $\nu=\nu_{s-1}$, $\bar{y}=\bar{y}^{s-2},$ $\tilde{l}=f\left(\bar{y}^{s-2}\right)-\Delta_{s-1},\tilde{y}=\bar{y}^{s-1}$, and $c=1$ yields that $\tilde{\mathcal{P}}_{s}=\left\{\bar{y}^{s-1}\right\}\cup\mathcal{P}_{s-1}$ is an $\left(\tilde{\iota}_{s},\tilde{\nu}_{s}\right)$-W-certificate for point $\bar{y}^{s-1}$ in function $f$, where $\tilde{\iota}_{s}=2\iota_{s-1}$ and
$\tilde{\nu}_{s}=\frac{1}{2\iota_{s-1}}\left(f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)+3\Delta_{s-1}\right)\ .$ (6.4)
@@
$f\left(\bar{y}^{s-1}\right)-f^{*}\leq\max\left\{\zeta_{s},\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2}\right\}\ .$ (6.5)
@@
$2\hat{\mu}_{s}\iota_{s-1}^{2}=\frac{4\hat{\mu}_{s}}{\mu_{s-1}}\Delta_{s-1}\overset{(a)}{\geq}\frac{4\hat{\mu}_{s}}{3\mu_{s-1}}\zeta_{s}\ ,$ (6.6)
where $\left(a\right)$ holds by $f\left(\bar{y}^{s-1}\right)-f\left(\bar{y}^{s-2}\right)\leq 0$. $\zeta_{s}>0$ and (6.6) imply $3\mu_{s-1}\zeta_{s}/4\hat{\mu}_{s}$ is an upper bound of the second term in (6.5). Note that $\hat{\mu}_{s}\leq\mu_{s-1}/2$, hence, we have
$\frac{3\mu_{s-1}}{4\hat{\mu}_{s}}\zeta_{s}\geq\max\left\{\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2},\frac{3}{2}\zeta_{s}\right\}\ .$ (6.7)
Hence, combining (6.7) and (6.5) yields
$\begin{split}f\left(\bar{y}^{s-1}\right)-f^{*}&\leq\max\left\{\zeta_{s},\frac{1}{2\hat{\mu}_{s}\iota_{s-1}^{2}}\zeta_{s}^{2},\frac{3}{2}\zeta_{s}\right\}\leq\frac{3\mu_{s-1}}{4\hat{\mu}_{s}}\zeta_{s}\end{split}\overset{(a)}{\leq}\frac{9\mu_{s-1}}{4\hat{\mu}_{s}}\Delta_{s-1}\ ,$ (6.8)